OpenStudy (anonymous):

-3(4b - 10) = 1 half(-24b + 60)

OpenStudy (anonymous):

step one distribute on the left

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

or is it not clear?

OpenStudy (anonymous):

if you want to impress your teacher you could alternatively divide both sides by \(-3\) as a first stop

OpenStudy (anonymous):

What about the 1 half.?

OpenStudy (anonymous):

oh i didn't see it did it just appear by magic or am i unable to read?

OpenStudy (anonymous):

nah, it wasn't there before you need to use the distributive law on both sides as the first step

OpenStudy (anonymous):

\[-3(4b-10)=\frac{1}{2}(-24b+60)\] right?

OpenStudy (anonymous):

Yeah thats it.

OpenStudy (anonymous):

remove parentheses using the distributive law first \[-12b+30=-12b+30\] is step one

OpenStudy (anonymous):

step two is to think "jeez, looks like the right hand side and left hand side are identical" so it must be an identity, i.e. true for all \(x\)